### v1

# Relative Overfitting and Accept-Reject Framework

Abstract
Currently, the scaling law of Large Language Models (LLMs) faces challenges and bottlenecks. This paper posits that noise effects, stemming from changes in the signal-to-noise ratio under diminishing marginal returns, are the root cause of these issues. To control this noise, we investigated the differences between models with performance advantages and disadvantages, introducing the concept of "relative overfitting." Based on their complementary strengths, we have proposed an application framework, Accept-Reject (AR). In Natural Language Processing (NLP), we use LLMs and Small Language Models (SLMs) as the medium for discussion. This framework enables SLMs to exert a universal positive influence on LLM decision outputs, rather than the intuitively expected negative influence. We validated our approach using self-built models based on mainstream architectures and pre-trained mainstream models across multiple datasets, including basic language modeling, long-context tasks, subject examination, and question-answering (QA) benchmarks. The results demonstrate that through our structure, compared to increasing the LLM’s parameters, we can achieve better performance improvements with significantly lower parameter and computational costs in many scenarios. These improvements are universal, stable, and effective. Furthermore, we explore the potential of "relative overfitting" and the AR framework in other machine learning domains, such as computer vision (CV) and AI for science. We hope the proposed approach can help scale laws overcome existing bottlenecks.
1 Introduction
Recently, the role of artificial intelligence in societal development has garnered increasing attention, particularly as the application of large language models (LLMs) significantly enhances human productivity. For an extended period, scaling laws 1 ; 2 ; 3 , driven by more data, greater computational power, and larger models, served as the guiding principle for training LLMs, allowing performance improvements through resource accumulation.
However, with the rapid advancement of artificial intelligence, scaling laws face increasing scrutiny 4 ; 5 , raising questions about whether they have reached a bottleneck. From the perspective of diminishing marginal returns, this observation is logical. While the debate continues over whether scaling laws have definitively hit a ceiling, evidence shows that the rate of performance improvement has notably decelerated. Furthermore, recent LLMs employing distillation techniques 6 ; 7 ; 8 ; 9 —where larger ’teacher’ models train smaller ’student’ models—demonstrate that SLMs can achieve performance comparable to LLMs (e.g., Deepseek R1 10 , Qwen 11 ). The DeepSeek R1 report specifically highlights that these distillation-based models with significantly fewer parameters perform remarkably close to their larger counterparts.
In this work, we explore an approach from the perspective of signal-to-noise ratio in model data, aiming to break the scaling law bottleneck. Data serves as the cornerstone for language models (LMs), yet along with its information, it inevitably introduces noise. Furthermore, the impact of the same data differs between a simple model exposed to limited data and a complex model utilizing vast amounts of data. Based on this premise, this paper introduces the concept of "relative overfitting" to better use the information and noise within a model. It posits that models with different structures exhibit fine-grained differences when modeling the target distribution. Within the same architecture, a model that better models the target distribution demonstrates relative overfitting compared to another model. Based on exploring this concept, we propose a universally applicable method.
Specifically, this paper proposes a universally applicable Accept-Reject (AR) framework and elaborates on it within the NLP domain, primarily using language models as vehicles. This choice is motivated by the abundance of scaling law-based model series in the field of natural language processing (NLP), which provides a robust experimental foundation. Although intuitively a model with inferior performance might seem to impede a superior one, particularly in the absence of sophisticated processing, we demonstrate that by adhering to specific proportions, a stable, positive enhancement in model performance can be achieved rather than a decline. This method, by balancing the relationship between model performance and relative overfitting, employs smaller models to influence larger ones through mechanisms of acceptance and rejection. Drawing upon the inevitable relative overfitting between models of different sizes, we utilize the Small Language Model (SLM) as a low-noise reference for the LLM. When the predictive discrepancy between them reaches a predefined threshold, we consider applying a corrective adjustment to the larger model. Compared to merely increasing the LLM’s parameter count, this structure, in many scenarios, can achieve equivalent or even superior performance enhancements for the LLM with significantly smaller parameter and computational costs. We have explored the threshold of the AR framework and posit that it could offer a potential avenue for addressing the bottlenecks in the scaling laws of large models. The method we propose is grounded in the internal relative overfitting issue within models and is not constrained by model architecture or absolute performance. It exhibits universal applicability across various model architectures and data benchmarks; that is, as long as the AR law is satisfied, even models with vastly different parameter magnitudes can be effectively influenced. Parameter magnitudes can be positively and effectively influenced.
We evaluated both smaller language models (SLMs) and larger language models (LLMs) across multiple datasets, comparing our custom-trained mainstream architectures with existing prominent language models. The results validate our method across several key aspects: its universality, shown by its applicability across all tested model architectures and datasets; in many scenarios, its effectiveness is shown by achieving better results with fewer parameters and computations, and its stability by the tendency for performance gains to increase with the main model’s size; and its flexibility, seen in offering considerably more adaptability compared to scaling up the LLM’s parameters or performing domain-specific fine-tuning. In addition, we also discussed the application of an AR framework based on the idea of relative overfitting in other areas of machine learning.
2 Related Work
The extent of actual parameter requirements has been previously investigated. Theoretically, based on the Johnson-Lindenstrauss lemma 13 , even a random projection, sampling a dimensionality reduction matrix from a Gaussian distribution, can compress N-dimensional features into $O(\log n/\epsilon^{2})$ dimensions. Experimentally, the findings of ALBERT 12 indicate that applying low-rank decomposition to BERT embeddings, reducing their dimensionality from 768 to 128, results in negligible performance degradation. This provides inspiration and evidence for our work, suggesting that when LM parameters are already substantial, there is parameter redundancy. Although increasing the parameters can enhance the granularity of the information, the associated cost of increased noise often outweighs this benefit. Furthermore, studies on word vectors, neural network interpretability, and LMs 14 ; 15 ; 16 ; 17 ; 18 ; 19 , together with research related to information and noise 20 ; 21 ; 22 ; 23 , have also informed our thinking.
Before this work, considerable research had been conducted on model fusion. Mainstream approaches include traditional ensemble methods, such as the multi-model voting schemes commonly used in Kaggle competitions 24 , which involve weighted combinations of outputs. Recent attempts within the LLM domain involve merging multiple models into a single entity 25 . Furthermore, some earlier studies have explored the fusion of output levels for LLM 26 . Unlike these studies, this article primarily explores the concept of relative overfitting and its derived application, the AR framework. Blending of models serves as a straightforward and comprehensible vehicle to articulate our ideas and validate the effectiveness of the proposed framework.
3 Relative Overfitting: What problems exist in LLM?
This chapter will utilize SLMs and LLMs as exemplars exhibiting relative overfitting (assuming that LLMs generally achieve higher granularity in modeling the target distribution than SLMs within the same architecture) in the NLP domain to elaborate on the theoretical foundation of our method: the concept of relative overfitting.
According to the scaling law of LMs, models with more parameters generally demonstrate a greater capacity to represent details. This essentially involves a trade-off between overfitting and underfitting. For models already proven to be excellent, both overfitting and underfitting are considered well controlled. Therefore, our focus is on "relative overfitting and underfitting." A larger model possesses higher semantic granularity compared to a smaller one; that is, an LLM exhibits relative overfitting compared to an SLM. This characteristic simultaneously drives its superior performance and creates limitations in other aspects. The underlying principle is the degree of fit to the target distribution. A model that fits the target distribution well inherently exhibits relative overfitting compared to a model that fits it poorly. Even with meticulous control, relative overfitting inevitably occurs compared to a lower-performing model. Consequently, this paper explores how to influence underfitting to influence overfitting, using LLMs and SLMs merely as vehicles for discussion, given that LLMs typically demonstrate relative overfitting concerning SLMs.
It is crucial to note that relative overfitting is not merely an extension of the concept of overfitting. Generally, overfitting refers to a scenario where model complexity, coupled with a small dataset, leads to excessive fitting of sample data, resulting in poor generalization to unseen datasets. However, in practice, well-tuned models following scaling laws, including LLMs, generally outperform SLMs even on unseen data. Our discussion centers on the fact that complex models possess a more detailed semantic space. Although both phenomena are due to model complexity, overfitting has associated control variables, and it can be managed by adjusting the training dataset without altering model parameters. However, relative overfitting lacks such an external control variable, making it impossible to manage this phenomenon by manipulating another quantity. This is why we emphasize the endogenous nature of relative overfitting. Simply put, overfitting depends on the relationship between model complexity and dataset size, whereas relative overfitting primarily depends on the model’s capacity to capture fine details of the target distribution.
In summary, relative overfitting enables one model to identify more semantic details than another. This means that LLMs can output sentences with substantial detail. Although generally beneficial, certain drawbacks that stem from this capability inspire our method. The issue arising from this higher semantic granularity is that although the output is more detailed, there is a greater likelihood of deviating from the core subject matter. LLMs strive to match the details of the target distribution as closely as possible, which consequently increases the probability of diverging from the main topic. Conversely, SLMs, constrained by their representational capacity, cannot inherently represent extensive detail and tend to concentrate on high-frequency words. In other words, an SLM might fail to represent adjectives, focusing instead on the high co-occurrence frequency of core components, whereas an LLM will pursue the most contextually appropriate and complete expression possible. Research indicates that in most scenarios, word distributions follow Zipf’s law. Furthermore, high-frequency words are preferentially fitted within deep neural network frameworks 53 ; 54 . The improvements offered by LLMs over smaller-parameter SLMs of the same architecture are primarily concentrated on low-frequency words; the fitting difference for extremely high-frequency words is relatively small between the two. Moreover, studies suggest that within commonly used frameworks, high-frequency words tend to be overestimated, while inaccurately fitted low-frequency words are generally underestimated 55 ; 56 ; 57 ; 58 .
We define $w_{1},w_{2},\dots,w_{V}\in V$ as word vectors sorted in descending order according to the true distribution $p(w|c)$, where $V$ represents the vocabulary and $c$ is the context set for prediction.
Based on established research findings and our experiments (detailed in Appendix B.1), we propose three hypotheses:
  • Words follow Zipf’s law distribution.
  • High-frequency words are preferentially fitted during training, and the fit error increases with decreasing frequency.
  • Based on experiments (Appendix B.1), we assume that the predicted probabilities $q(w|c)$ satisfy: $q_{\text{SLM}}(w_{k}|c)\approx 0$ for $k>i$, while $q_{\text{LLM}}(w_{k}|c)\approx 0$ primarily for $k>j$, with $q_{\text{LLM}}(w_{k}|c)>0$ for some $i<k\leq j$.
We define the "core" as the high-frequency region $w_{1}$-$w_{i}$. This leads to our key theoretical result: Theorem 1. Under the proposed hypotheses, let $a\in[0,i]$, $b\in[i+1,j]$. The probability of an LLM deviating from the core region exceeds that of an SLM, formalized as:
$\frac{\sum_{a=0}^{i}|p(w_{a}|c)-q_{LLM}(w_{a}|c)|+\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{LLM}(w_{b}|c)|}{j}\geq\frac{\sum_{a=0}^{i}|p(w_{a}|c)-q_{SLM}(w_{a}|c)|}{i}$ (1)
The complete formalization of hypotheses and proofs is provided in Appendix A.
Based on the ideas discussed above, it can be articulated outside the field of NLP that there exists a relative overfitting of the larger side in scaling laws compared to the smaller side. This leads to a better modeling capability of the details of the target distribution but also increases the likelihood of deviating from the ’core’.
4 AR Framework
Based on the discussion of relative overfitting, we leverage this property to propose the ’AR framework.’ This approach utilizes the strengths of the model with a comparatively weaker fitting capability for the target distribution (in the context of relative overfitting) to influence the other model. Without loss of generality, we will continue to use SLMs and LLMs as representative models for our discussion in the NLP domain. The entire model will be introduced and discussed in four sections. In Section 4.1, we will introduce a simple application method based on the AR framework. Section 4.2 introduces the fundamental principles proposed based on the AR framework. In Section 4.3, we discuss the optimal threshold point for this framework. Finally, Section 4.4 discusses the magnitude of performance improvement achieved by the model.
4.1 Our Method
As demonstrated in the previous chapter, LLMs of the same architecture exhibit a higher probability of deviating from the core path compared to SLMs. We propose an AR framework based on this relative overfitting, specifically, a method using the SLM to influence the LLM. The underlying principle is that we balance the SLM and LLM based on the concept of relative overfitting to leverage their respective advantages, ensuring that in most situations, the SLM does not adversely affect the LLM, meaning its judgments are accepted. However, in extreme cases, the SLM can exert a rectifying effect on the LLM, leading to the rejection of the LLM’s judgment. The rationale behind this is that the existence of relative overfitting makes the LLM more prone to deviating from the core path. Therefore, if the LLM’s output significantly diverges from the SLM’s judgment, we infer that it has strayed from the correct core path. This approach allows us to preserve the LLM’s advantage in modeling detailed aspects of the target distribution under normal circumstances while exploiting the SLM’s superior advantage in modeling the core distribution (particularly concerning high-frequency words) in exceptional cases.
Based on the AR framework, we introduce a simple application training-free method. The specific implementation is as follows: utilizing our proposed AR law, we allocate proportions between the SLM and LLM, thereby identifying the AR’s threshold point. Based on this threshold point, we assign weights to the smaller and larger models and perform a weighted combination at the probability distribution layer of their outputs to obtain the final result. The weights assigned based on this threshold point (acting as a fulcrum) effectively establish a threshold to determine significant deviation. For the LLM’s output, the SLM, holding minimal weight, generally has no impact, meaning that in most situations, we primarily utilize the LLM’s detailed advantage and accept its judgment. However, once the conflict between the two reaches a critical point, that is, the SLM strongly contradicts the LLM’s judgment, we have reason to believe the LLM’s output is progressively deviating from the core path. At this juncture, we consider rejecting the LLM’s judgment. This enables us, in extreme situations, to take advantage of the core advantage of the SLM, thus achieving complementary advantages between the models (as illustrated in Figure 1).
Figure 1: A simple manifestation of an AR framework.
It is important to note that our emphasized viewpoint is that such discrepancies are an endogenous issue that inevitably arises from relative overfitting caused by the model structure. That is, under general circumstances, even LMs with vast performance differences possess their respective strengths and weaknesses. Our AR law allows for the effective utilization of these advantages, however slight they may seem. This forms the fundamental basis for the general and stable efficacy of our method. Naturally, in extreme cases, such as a model with exceedingly poor performance, the core region we define approximates an empty set, resulting in the SLM’s weight becoming zero, thereby exerting no influence. Specifically, applying softmax to the logit vectors $l_{i}$ ($i=1,2$) output by the two models yields:
$p_{i}=\mathrm{softmax}(l_{i})$ (2)
Based on the threshold point of AR, an appropriate weight $\alpha$ is determined, and a weighted sum is performed to obtain the final output:
$p_{\text{ensemble}}=\alpha\cdot p_{1}+(1-\alpha)\cdot p_{2}$ (3)
It can be asserted that an SLM satisfying the threshold point of the AR condition will invariably exert a positive influence on the LLM. This influence is determined by the relative overfitting between the different models, is endogenous, and represents an inevitable outcome of the trade-offs involved. This explains its universal applicability, irrespective of model architecture or performance level. We discuss this point further in the subsequent experimental section, noting that while the magnitude of performance improvement may vary, an enhancement is consistently achievable.
Currently, when considering parallel implementation, this method incurs only linear complexity related to the vocabulary size. In many situations, compared to improving LLM performance by increasing the number of parameters, this approach offers significant savings in memory and computation time.
4.2 AR Law
The AR law constitutes the rule within the AR framework that defines the acceptance and rejection criteria for judgments of various SLMs when operating with a fixed LLM. This chapter will illustrate this based on the ideas, discussions, and experimental results of our approach, using the simple implementation under the AR framework described in Section 4.1 as an example.
First, we introduce the concept of the threshold point of AR. Let the parameter ratio between the LLM and the SLM be $k=\frac{|\theta_{\mathrm{LLM}}|}{|\theta_{\mathrm{SLM}}|}$, where $|\theta_{\mathrm{LLM}}|$ denotes the number of parameters in the LLM, and $|\theta_{\mathrm{SLM}}|$ denotes the number of parameters in the SLM. Since the performance of the same model can vary across different datasets, for convenience, this paper uses parameter count as a proxy for the concept of performance. It is important to note that although our discussion uses the parameter ratio as a metric, it represents the performance ratio and the underlying concept of relative overfitting. Therefore, if the AR weight allocation ratio that achieves the optimal performance improvement is $\alpha$, then we have:
$\alpha\propto k$ (4)
Second, we introduce the concept of the improvement domain, denoted as $[R,\infty]$, where $R<\alpha$. For any weight allocation ratio $\beta$, when $\beta\in[R,\infty]$, the model can achieve a performance improvement. Furthermore, when $\beta<\alpha$, the magnitude of performance improvement increases, and when $\beta>\alpha$, the magnitude of performance improvement decreases. When $\beta=\infty$, it approximates a single model. We also have:
$R\propto k$ (5)
When $k$ is extremely large, meaning the performance (parameter) gap between the models is vast, $R$ approaches infinity ($R\to\infty$). In this scenario, we consider the performance improvement to be near zero. Of course, this situation rarely occurs. To verify this extreme case, we designed an experiment using a model with extremely poor performance and very high perplexity to construct the AR framework. In this case, the performance improvement was almost negligible.
Finally, the magnitude of performance improvement is inversely proportional to the weight ratio $\beta$.
The points described above constitute the basic AR law defined on the basis of the theoretical ideas presented earlier. Next, we will discuss the selection of the threshold point of AR.
4.3 The Threshold Point of AR
Based on the ideas discussed in the previous section, we examine the optimal allocation ratio threshold point $\alpha$ based on experimental results. We consider two main influencing factors in the entire system:
  • Marginal effect of the performance gap between LLM and SLM: In most cases, we tend to leverage details, meaning that as the performance gap widens, the LLM should be assigned a greater weight.
  • Relative overfitting: Relative overfitting tends to shift the leverage weight ratio slightly towards the SLM.
We decompose the optimal ratio $\alpha$ into two conflicting components. The first is an ascending component. If the parameter ratio $k$ between the LLM and SLM is large, we should use a larger $\alpha$ to better utilize the LLM’s capability for detail, essentially, its conventionally understood performance. According to Kaplan 1 , LLM performance follows a power law concerning parameter count, suggesting($p$ is the power-law exponent):
$\alpha\propto k^{p}$ (6)
The second component characterizes the magnitude at which the AR framework operates effectively. In reality, $k$ only describes the gap between the LLM and SLM but fails to capture the starting point of this gap. We introduce the parameter $\theta_{LLM}$ (the LLM’s parameter count) to represent the starting point where this influence takes effect. For example, the ratio between parameters 14B, 7B, 14M, and 28M is also 2, but the magnitude of the differences is entirely different. The relative overfitting cases arising from different magnitudes will also vary. Therefore, as noted in Section 4.2, we are discussing how to influence it based on the AR framework and AR law under a fixed LLM. The average optimal weight $\alpha$ of each model across various datasets in all experiments from Section 5.1 is presented in Appendix B.2.
4.4 Improvement Magnitude
In this subsection, we discuss the magnitude of performance improvement contributed by the lower-performing model to the higher-performing model within the AR framework. Differing from previous discussions, the magnitude of performance improvement, while dependent on the weights, is influenced by more complex factors in practice than the relatively strict proportional and inverse relationships explored in Section 4.2. Based on experimental results, we identify two empirical regularities. First, as the weight ratio (associated with the LLM) increases, the proportion and influence of the SLM diminish. However, this inverse relationship is not absolute; exceptions were observed in experiments. This is because the change in proportion is an integral part of the AR law itself—reducing the SLM’s proportion serves the mechanism’s requirements and is not merely about decreasing the SLM’s influence. we see that the actual optimal weight ratio $\alpha$ increases only with $k$ (the parameter ratio).
Secondly, as stated in Section 4.3. We observe that while the actual optimal weight ratio $\alpha$ increases with our experiments indicate it does not demonstrably increase with the scale of the main model parameters $\theta_{LLM}$. This provides evidence for the stability of the AR framework: namely, as long as an appropriate k is maintained, the increase in the model’s performance remains stable with the growth of $\theta_{LLM}$, and in some scenarios, there is even a rising trend (as seen in Section 5.1). This means that, according to the AR law, we need not worry about the significant reduction in performance due to excessively large main model parameters. This finding supports our belief that this approach can positively impact the bottleneck observed in scaling laws. And under the scaling law, as the parameters of LLM increase, the rate of improvement decreases rapidly.
5 Experiments
To validate the method proposed in this paper, in the field of NLP, we designed experiments considering existing mainstream scaling-law-based LMs as well as custom-built LMs based on prevalent architectures. The aim is to verify the universality, effectiveness, stability, and flexibility of our approach. First, we tested our method on several existing scaling-law-based model series (e.g., GPT-2 33 ). Second, we conducted ablation studies using a model built upon the Transformer 31 architecture, deliberately avoiding any task-specific design modifications, and comparing against existing models. Finally, we retrained models of varying scales based on the mainstream architectures LSTM 29 ; 30 , Transformer, and GRU 32 for further experimentation. In summary, this chapter demonstrates that the proposed method is universally applicable across a variety of models, including mainstream models, custom-built models, and combinations thereof. Furthermore, we also validated the AR framework, based on the principle of relative overfitting, in the domains of computer vision (CV) and AI for science.
5.1 Mainstream Language Model Testing
This paper conducted experiments using the proposed AR framework on several pre-trained models. Tests were performed using GPT-2 33 on fundamental language modeling benchmarks including WikiText2, WikiText103 34 , text8 14 , Pile 39 ; 40 (using a randomly sampled subset of the Pile test set), and 1BW 38 , as well as on the long-context benchmark LAMBADA 35 and QA benchmarks CBT 36 and subject examination benchmark ARC 37 . We also evaluated the performance of models with different architectures—Pythia 41 (multi-scale scaling law), RWKV 42 (combining RNN and Transformer), Mamba 43 (SSM series), and Qwen 44 ; 45 (commercial LLM)—using the AR framework for basic language modeling tasks. To compare the behavior under the AR framework between general language modeling and specific tasks, we used the LAMBADA dataset for both language modeling evaluation (perplexity) and its original long-context dependency evaluation.
The results in Table 1 demonstrate that through our AR framework, all models achieve universal performance improvements. In many scenarios, we can achieve superior performance with significantly fewer parameters and less computation time. For instance, on LAMBADA (PPL), Pile, and other benchmarks, our method consistently outperforms the strategy of simply increasing the LLM’s parameter count, achieving superiority in terms of parameters, computation time, and performance. This advantage becomes more pronounced as the model parameters increase. For example, based on the improvement rate observed when scaling GPT-2 from 762M to 1.5B, and roughly estimating the improvement from 1.5B to 3B using the power-law model proposed in scaling laws, the performance of the 1.5B+762M combination almost universally surpasses the projected 3B model across all benchmarks. Notably, our method’s advantages become more significant with larger models, while the linear complexity related to vocabulary size becomes negligible compared to increasing the parameter count. Furthermore, as the main model’s parameters grow, the magnitude of performance improvement increases rather than diminishes, leading us to be optimistic about its performance on even larger models and hopeful that it can contribute to alleviating the scaling law bottleneck.
Table 1: The table shows the test results of the classic model GPT2 based on the AR framework with four parameter sizes from 117M to 1.5B. LTD is the long context benchmark under the LAMBADA dataset. The results of RWKV, Mamba, Pythia, and Qwen can be found in the appendix.
LLM SLM WikiText2 WikiText103 text8 1BW Pile(part) LAMBADA LAMBADA CBT-CN CBT-NE ARC
GPT2 GPT2 PPL $\downarrow$ PPL $\downarrow$ BPC $\downarrow$ PPL $\downarrow$ PLL $\downarrow$ PPL $\downarrow$ PLL(LTD) $\downarrow$ ACC $\uparrow$ ACC $\uparrow$ ACC $\uparrow$
117m 23.52 29.16 1.21 63.97 18.66 43.08 24.99 78.84 72.40 22.35
345m 17.99 21.08 1.10 52.75 13.74 33.66 12.66 82.88 75.36 25.00
762m 15.84 18.15 1.05 45.67 13.20 30.61 9.33 84.68 76.32 24.40
1.5b 14.80 16.51 1.02 43.81 12.15 29.96 7.80 86.92 78.92 26.62
Improvement of the evaluation metric relative to the best performing single model (%)
345m 117m 1.15% 0.62% 0.32% 2.16% 1.11% 1.63% 0.28% 0.24% 2.34% 1.71%
762m 117m 0.77% 0.42% 0.20% 1.15% 4.68% 1.81% 0.01% 0.43% 2.46% 1.40%
762m 345m 2.73% 2.17% 1.00% 2.26% 9.51% 4.13% 1.07% 0.85% 1.52% 3.41%
1.5b 117m 0.82% 0.50% 0.27% 1.52% 4.52% 2.36% 0.12% 0.37% 1.57% 1.28%
1.5b 345m 2.01% 1.47% 0.72% 2.16% 8.41% 4.92% 0.54% 0.18% 0.66% 2.88%
1.5b 762m 3.85% 3.24% 1.38% 4.72% 4.97% 6.99% 3.31% 0.18% 0.61% 1.92%
5.2 Ablation Experiments with Custom-Built Models Based on Mainstream Architectures
To verify the universality and flexibility of our proposed method, in this section, we utilized the vocabularies corresponding to Pythia, RWKV, Mamba, and GPT-2. Without incorporating any specific design elements, we employed a simple Transformer architecture to train a reference model with only approximately 180M parameters. In practice, the standalone performance of this model is not particularly strong.
To further validate the universality, stability, and flexibility of our method, we designed a control experiment. On the WikiText103 and LAMBADA datasets, we performed predictions using only the model trained on WikiText2, without using the corresponding training sets for these target datasets. This was done to confirm that the performance improvement achieved by the AR framework is not merely due to having "seen the data". Additionally, to verify the effectiveness of our method, we trained models on the 1BW and text8 datasets using their respective training sets, aiming to achieve substantial performance gains.
The results in Table 2 indicate that although the models we trained on WikiText2 and WikiText103 did not perform well individually, their improvement magnitudes exceeded those of inherently superior but highly homogeneous models. The results on the LAMBADA dataset further demonstrate that even extremely poor-performing models can still contribute improvements through the AR framework. The outcomes on the text8 and 1BW datasets illustrate that our method is not only universal but also highly effective (even though we did not specifically search for the optimal fulcrum $\alpha$). The AR framework based on non-homogeneous models has nearly achieved comprehensive advantages in computation, parameters, and performance in certain scenarios compared to LLMs that seek improvement by increasing the number of parameters. Even without additional design, enhancements will occur as long as an appropriate threshold point is identified based on relative overfitting using the AR framework.
It is important to note that this is not fine-tuning 46 ; 47 , but merely a training-free bias correction. Compared to fine-tuning, our method offers greater flexibility, requires significantly fewer resources, and does not affect the base model (domain-specific fine-tuning may lead to performance degradation in other domains, as seen with DeepSeekR1 48 ).
Table 2: All models used the simplest Transformer architecture. For WikiText2, WikiText103, and LAMBADA results, models were trained on the WikiText2 training set, while for 1BW and text8 results, models were trained on the 1BW and text8 training sets, respectively.
Model WikiText2 WikiText103 LAMBADA text8 1BW
Trained Model PPL $\downarrow$ PPL $\downarrow$ PPL $\downarrow$ BPC $\downarrow$ PPL $\downarrow$
GPT2-Train 91 322 9084 1.03 131
Other-Train 94 314 12084 1.10 45
Improvement of the evaluation metric relative to the best performing single model (%)
GPT2-1.5b 9.46% 6.33% 1.02% 26.12% 22.40%
Pythia-2.8b 3.76% 3.09% 0.10% 19.61% 47.74%
Mamba-2.8b 4.26% 3.56% 0.10% 18.68% 45.71%
RWKV-1.5b 4.64% 3.67% 0.08% 20.53% 49.02%
Figure 2: Test results of AR framework based on simple LSTM, Transformer and GRU architectures under different embedding dimensions of SLM and LLM..
5.3 Self-built model testing based on mainstream architectures
This paper conducted tests based on the current mainstream model architectures LSTM 29 ; 30 , Transformer 31 , and GRU 32 , utilizing the same vocabulary but varying embedding dimensions. In practice, we employed very simple architectures 49 ; 50 ; 51 ; 52 without any specific model design modifications. Even under these conditions, our method continues to demonstrate universal and stable effectiveness (as illustrated in Figure 2).
5.4 Experiments in the rest of the machine learning field
Based on the principles of relative overfitting and the AR framework, we employed a similar methodology to conduct experiments with our approach across a broader range of machine learning domains, including CV and AI for Science, thereby validating the general effectiveness of our proposed framework (further details are available in Appendix C).
6 Discussion and Future work
We discuss limitations and future directions.
  • Model Selection Trade-offs: For homogeneous scaling models, while performance showed stability (no significant fluctuations) as the main model grew, the improvement magnitude diminished as the parameter ratio expanded. Larger ratios necessitated more extreme leverage, reducing SLM effectiveness and potentially limiting resource savings. Thus, in specific domains, non-homogeneous models might be a better choice.
  • Better Improvement Magnitude: Although this paper proposes the AR law based on relative overfitting and demonstrates its universal, stable, and effective improvements along with good flexibility under a suitable threshold point, significant room for progress remains beyond the decent performance improvements already achieved. To reveal fundamental principles, our method attained these results despite lacking any specific model design, highlighting the potential yet to be unlocked. Therefore, with a deeper understanding of the entire system and relative overfitting, key research questions worth exploring include different combination methods based on the AR law, the selection and innovation of low-parameter reference model architectures serving as backbones, and how to better maintain the information purity of the SLM. We anticipate that addressing these questions will lead to the emergence of greater improvement magnitudes.
  • Limitations of the AR Law: As noted in this paper, the AR framework involves complex relative overfitting. Although we have proposed fundamental principles applicable in most scenarios, the inherent black-box nature and uncertainty of neural networks mean that minor fluctuations can occur within the general direction established by the AR Law. Consequently, we believe that the complex influence mechanisms of relative overfitting warrant further in-depth research.
  • Broader Applicability: Although this paper primarily focuses on the NLP domain and has also conducted some experiments in CV and AI for Science, due to limited time and resources, a broader discussion across more fields was not feasible. The applicability of the relative overfitting phenomenon and the proposed AR framework in a wider range of domains warrants further exploration.
Appendix
A Proof of Theorem 1
Let $V$ be the vocabulary size. We define $w_{1},w_{2},\dots,w_{V}\in V$ as word vectors sorted in descending order according to the true distribution $p(w|c)$, where $c$ represents the context set for prediction, $c\in C$. LLM and SLM refer to models based on similar architectures but differing in complexity.
The vocabulary $W$ is partitioned into three disjoint regions based on word frequency ranks derived from the true distribution $p(w|c)$:
  • High-frequency region (Core): $W_{core}=\{w_{1},\dots,w_{i}\}$(Index 1 to i as per definition)*
  • Mid-frequency region: $W_{mid}=\{w_{i+1},\dots,w_{j}\}$
  • Low-frequency region: $W_{low}=\{w_{j+1},\dots,w_{V}\}$
where $1\leq i<j<V$. The indices $i$ and $j$ define the boundaries based on observed model capabilities and frequency distinctions.
Based on existing research findings, we propose the following reasonable hypotheses:
  • Hypothesis 1:(Zipf’s Law) The distribution of words follows Zipf’s law.
  • Hypothesis 2:(Fitting Trend and Core Consistency) (a) (Error Trend) During language model training, high-frequency words are preferentially fitted. The absolute fitting error $|p(w_{k}|c)-q(w_{k}|c)|$ generally increases as rank $k$ increases.
    This implies that the average fitting error in the mid-frequency region $W_{mid}$ is higher than in the core region $W_{core}$.
    Let $E_{core}^{M}=\sum_{a=1}^{i}|p(w_{a}|c)-q_{M}(w_{a}|c)|$ denote the total core error, and $E_{mid}^{M}=\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{M}(w_{b}|c)|$ denote the total mid-region error. Define the average errors (for $j>i$) as:
    $\bar{e}_{core}^{M}=\frac{E_{core}^{M}}{i}\quad\text{and}\quad\bar{e}_{mid}^{M}=\frac{E_{mid}^{M}}{j-i}.$ (7)
    The implication of the error trend is $\bar{e}_{mid}^{LLM}>\bar{e}_{core}^{LLM}$.
    (b) (Core Consistency) The fitting capabilities of LLM and SLM in the high-frequency core region $W_{core}$ (ranks 1 to i) are consistent, meaning their average errors are approximately equal:
    $|\bar{e}_{core}^{LLM}-\bar{e}_{core}^{SLM}|<\epsilon$ (8)
    where $\epsilon$ is a small positive constant.
  • Hypothesis 3:(Model Output Behavior)
    (1) Both models predict effectively in $W_{core}$ ($k\in[1,i]$). *(Redundant given H2(b), kept for clarity of structure)*
    (2) In $W_{mid}$ ($k\in[i+1,j]$), $q_{LLM}(w_{k}|c)>0$ for all $k$, while $q_{SLM}(w_{k}|c)\approx 0$.
    (3) In $W_{low}$ ($k\in[j+1,V]$), $q_{LLM}(w_{k}|c)\approx 0$ and $q_{SLM}(w_{k}|c)\approx 0$.
Based on the above hypotheses, we define the core as the high-frequency region $w_{1}$-$w_{i}$ and obtain the following theoretical result regarding the model’s average fitting error:
Let $a\in[1,i]$, $b\in[i+1,j]$. The average fitting error of LLM over the range [1, j] is greater than or equal to that of SLM over the range [1, i]:
$\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{LLM}(w_{a}|c)|+\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{LLM}(w_{b}|c)|}{j}\geq\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{SLM}(w_{a}|c)|}{i}$ (9)
Let $LHS=\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{LLM}(w_{a}|c)|+\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{LLM}(w_{b}|c)|}{j}$. Let $RHS=\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{SLM}(w_{a}|c)|}{i}$. We aim to prove $LHS\geq RHS$.
Define the average errors as in Hypothesis 2: $\bar{e}_{core}^{M}=\frac{\sum_{a=1}^{i}|p(w_{a}|c)-q_{M}(w_{a}|c)|}{i}$ for the core region [1, i]. $\bar{e}_{mid}^{LLM}=\frac{\sum_{b=i+1}^{j}|p(w_{b}|c)-q_{LLM}(w_{b}|c)|}{j-i}$ for the mid region [i+1, j] (assuming j>i).
From Hypothesis 2(b) (Core Consistency), we have:
$|\bar{e}_{core}^{LLM}-\bar{e}_{core}^{SLM}|<\epsilon$ (10)
So, $RHS=\bar{e}_{core}^{SLM}\approx\bar{e}_{core}^{LLM}$.
From Hypothesis 2(a) (Error Trend), we have:
$\bar{e}_{mid}^{LLM}>\bar{e}_{core}^{LLM}\quad(2)$ (11)
Now, express LHS in terms of average errors:
$LHS=\frac{i\cdot\bar{e}_{core}^{LLM}+(j-i)\bar{e}_{mid}^{LLM}}{j}$ (12)
Substitute $\bar{e}_{mid}^{LLM}$ using inequality (2). Since $\bar{e}_{mid}^{LLM}>\bar{e}_{core}^{LLM}$, we replace $\bar{e}_{mid}^{LLM}$ with the smaller value $\bar{e}_{core}^{LLM}$ to obtain a strict inequality:
$LHS>\frac{i\cdot\bar{e}_{core}^{LLM}+(j-i)\bar{e}_{core}^{LLM}}{j}=\frac{(i+j-i)\bar{e}_{core}^{LLM}}{j}=\frac{j\bar{e}_{core}^{LLM}}{j}=\bar{e}_{core}^{LLM}$ (13)
So, we have shown:
$LHS>\bar{e}_{core}^{LLM}\quad(3)$ (14)
Combining (1) and (3):
$LHS>\bar{e}_{core}^{LLM}\approx\bar{e}_{core}^{SLM}=RHS$ (15)
Since $LHS$ is strictly greater than a value approximately equal to $RHS$, the non-strict inequality $LHS\geq RHS$ holds.
Therefore, the theorem is proven.
B Results of Chapters Three and Four Experiments
B.1 Larger models have more optional tokens
Figure 3: Proportion of tokens generated by scaling law model with different parameters to the total tokens of the dataset under different datasets.
The number of tokens that the model may generate across all datasets increases with the increase in model parameters.
B.2 Larger models have more optional tokens
Figure 4: Variation of Optimal Weight Ratio $\alpha$ with Parameter Ratio $k$ for Fixed Main Parameter $\theta_{LLM}$.
From Figure 4, it can be observed that under the condition of the same main $\theta_{LLM}$ parameters, the optimal weight ratio $\alpha$ for each model shows an increasing trend with the rise in the parameter ratio.
C Experiments in the rest of the machine learning field
In this appendix, we detail the experimental setups employed. For image recognition, a suite of pre-trained ResNet models of varying sizes (ResNet-18 to ResNet-152) 64 was utilized. These models were subsequently fine-tuned for one epoch on the CIFAR-10, CIFAR-10067 , and Tiny ImageNet-200 datasets 59 ; 60 ; 61 ; 62 ; 63 to assess classification efficacy across varying complexity levels. Concurrently, for the protein sequence prediction task, the ESM-265 model was employed. This task focuses on predicting protein sequences, which holds significant societal importance, including applications such as the design of novel protein sequences for therapeutic or industrial purposes. The model’s predictive capabilities were benchmarked on the UniRef5066 dataset.
The experimental results demonstrate that the ideas and framework we have proposed can be universally applied to other areas of machine learning.
C.1 Computer Vision
In our experiments, we employed a consistent post-training approach across CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. We fine-tuned ResNet series models (from ResNet18 to ResNet152) pre-trained on ImageNet. The fine-tuning process began by replacing the final fully connected layer, adjusting the output dimensions to match the respective datasets: 10 classes for CIFAR-10, 100 for CIFAR-100, and 200 for Tiny-ImageNet. All images were resized to 224×224 pixels and normalized using ImageNet parameters (mean=[0.485, 0.456, 0.406], std=[0.229, 0.224, 0.225]). We utilized SGD optimizer with a learning rate of 0.001 and momentum of 0.9, typically running for one epoch. Performance was evaluated using Top-1 accuracy metrics for both individual models and ensembles.
Table 3: Top-1 Accuracy of ResNet models and AR framework ensembles on CIFAR-10, CIFAR-100, and Tiny ImageNet. Accuracies (%) and relative improvements (%) are shown. Improvements are calculated for the Large Models (LM) when combined with the Small Models (SM), relative to the best-performing single model’s Top-1 performance.
Model Configuration CIFAR-10 CIFAR-100 Tiny ImageNet
Single Model Baseline Performance (Top-1 Acc. (%))
resnet18 90.44 55.47 38.69
resnet34 93.14 64.99 49.66
resnet50 93.69 65.54 55.19
resnet101 95.31 71.46 64.61
resnet152 95.39 72.80 67.18
AR Framework Ensemble: Top-1 Improvement relative to LM Baseline (%)
LM SM Top-1 Imp. (%) ↑
resnet34 resnet18 0.61% 1.55% 2.07%
resnet50 resnet18 0.21% 1.19% 0.49%
resnet50 resnet34 0.81% 5.00% 3.26%
resnet101 resnet18 0.17% 0.57% 0.01%
resnet101 resnet34 0.41% 1.43% 0.12%
resnet101 resnet50 0.39% 1.33% 0.51%
resnet152 resnet18 0.52% 0.84% 0.01%
resnet152 resnet34 0.29% 1.70% 0.06%
resnet152 resnet50 0.43% 1.48% 0.13%
resnet152 resnet101 0.85% 2.90% 2.22%
C.2 AI For Science
We evaluated different scales of ESM-2 protein language models (8M to 3B parameters) using the Masked Language Modeling (MLM) task. The experiment employed a 15% random masking rate on amino acid residues while avoiding special tokens ([CLS], [SEP], [PAD], etc.). Performance was measured using perplexity (PPL), calculated as the exponentiated cross-entropy loss on masked positions. For excessively large loss values (>70), we returned infinity to prevent numerical overflow. This evaluation methodology effectively reflects the protein sequence modeling capabilities of ESM-2 models across different parameter scales.
Table 4: Perplexity (PPL) of ESM-2 models and AR framework ensembles on UniRef50 (test_00.csv). Baseline PPLs and relative PPL reductions (%) for ensembles are shown. PPL reductions are calculated for the LM when combined with the SM, relative to the best performing single baseline PPL.
Model Configuration PPL $\downarrow$
Single Model Baseline Performance (PPL)
esm2_8m 12.30
esm2_35m 10.65
esm2_150m 9.17
esm2_650m 7.59
esm2_3b 6.65
AR Framework Ensemble: PPL Reduction relative to LM Baseline (%)
LM SM PPL Reduction (%) $\uparrow$
esm2_35m esm2_8m 0.56%
esm2_150m esm2_8m 0.12%
esm2_150m esm2_35m 0.36%
esm2_650m esm2_8m 0.17%
esm2_650m esm2_35m 0.14%
esm2_650m esm2_150m 0.01%
esm2_3b esm2_8m 0.17%
esm2_3b esm2_35m 0.17%
esm2_3b esm2_150m 0.18%
esm2_3b esm2_650m 0.78%
The experimental results of using ResNet for image classification and ESM2 for protein sequence prediction demonstrate that the straightforward method proposed based on the AR framework is still generally effective in fields beyond NLP. However, it is important to note that in non-language modeling problems, the performance of the AR law does not exhibit the same strictness as it does in language modeling problems. While the overall trend aligns with the AR law, small fluctuations may occur locally. This is because the AR framework is based on relative overfitting for correction, which is quite intuitive in basic language modeling problems, but exhibits more complexity in other scenarios. Therefore, this paper posits that further in-depth research on relative overfitting and the AR law is necessary.
D Complementary Experiments
In Appendix D, we demonstrate the performance of Pythia, RWKV, Mamba, and Qwen (based on int4 quantization) within the Qwen framework, which is based on the AR framework, on the fundamental language modeling benchmarks WikiText2, WikiText103, text8, and Pile (with Pile being a randomly sampled portion of the test set), as well as the 1BW benchmark. We found that the improvements under our framework are generally effective and largely conform to the AR law we proposed.
Qwen is a series of large language models developed by Alibaba Cloud in 2023. As a prominent commercial large language model, it competes with other industry models like GPT and Claude. The model architecture is based on the Transformer decoder framework, with optimizations for context length and instruction-following capabilities. We used five different sizes of the Qwen model: 0.5b, 1.5b, 3b, 7b, and 14b. To facilitate formatting, we will present the quantization results based on int4 in Table 8.
Table 5: The table lists the test results of the commercial LLM Qwen based on our proposed AR framework in fundamental language benchmarks.
LLM SLM WikiText2 WikiText103 text8 1BW Pile(part) LAMBADA
Qwen Qwen PPL $\downarrow$ PPL $\downarrow$ PPL $\downarrow$ BPC $\downarrow$ PLL $\downarrow$ PPL $\downarrow$
Single Model Baseline Performance
0.5b 12.41 14.56 1.00 40.87 10.74 37.22
1.5b 9.48 10.49 0.86 30.66 8.15 29.17
3b 8.51 9.19 0.80 26.79 7.30 26.23
Improvement of the evaluation metric relative to the best performing single model (%)
1.5b 0.5b 1.47% 1.42% 0.37% 1.21% 0.46% 1.21%
3b 0.5b 1.10% 1.15% 0.20% 0.66% 0.21% 0.65%
3b 1.5b 3.18% 3.29% 1.17% 2.50% 1.35% 2.27%
Receptance Weighted Key Value (RWKV) is a novel architecture that combines the parallelizable training of Transformers with the constant-memory, linear-time inference of RNNs by leveraging a linear attention mechanism, achieving performance on par with similarly sized Transformers. In our experiments, we used the 169M, 430M, and 1.5B variants of RWKV.
Table 6: The table lists the test results of RWKV, a model that fuses RNN and Transformer structures, based on our proposed AR framework in fundamental language benchmarks.
LLM SLM WikiText2 WikiText103 text8 1BW Pile(part) LAMBADA
RWKV RWKV PPL $\downarrow$ PPL $\downarrow$ PPL $\downarrow$ BPC $\downarrow$ PLL $\downarrow$ PPL $\downarrow$
Single Model Baseline Performance
169m 24.10 29.19 1.23 56.70 13.31 34.91
430m 17.23 19.43 1.09 43.86 10.10 28.59
1.5b 12.76 13.71 0.97 33.39 8.05 23.93
Improvement of the evaluation metric relative to the best performing single model (%)
430m 169m 1.88% 1.11% 0.40% 1.70% 0.65% 1.31%
1.5b 169m 0.26% 0.17% 0.10% 0.34% 0.28% 0.48%
1.5b 430m 1.18% 0.74% 0.38% 1.00% 0.62% 1.04%
Mamba is a linear-time sequence modeling architecture that replaces traditional attention modules with input-conditioned Selective State Space Model (SSM) layers, enabling content-based reasoning over long sequences in O(n) time while maintaining Transformer-comparable accuracy. Mamba has demonstrated representative performance across modalities such as language, audio, and genomics. In our experiments, we used the 130M, 370M, 790M, 1.4B, and 2.8B variants of Mamba.
Table 7: The table lists the test results of Mamba, a model using the SSM architecture, based on our proposed AR framework in fundamental language benchmarks.
LLM SLM WikiText2 WikiText103 text8 1BW Pile(part) LAMBADA
Mamba Mamba PPL $\downarrow$ PPL $\downarrow$ PPL $\downarrow$ BPC $\downarrow$ PLL $\downarrow$ PPL $\downarrow$
Single Model Baseline Performance
130m 19.59 23.26 1.14 51.04 11.70 31.71
370m 14.40 16.05 1.02 38.55 9.05 26.13
790m 12.49 13.51 0.95 32.99 7.95 23.51
1.4b 11.44 12.16 0.92 30.58 7.40 22.32
2.8b 10.21 10.60 0.87 27.95 6.69 20.69
Improvement of the evaluation metric relative to the best performing single model (%)
370m 130m 1.33% 0.83% 0.38% 1.00% 0.52% 1.26%
790m 130m 0.68% 0.35% 0.20% 0.39% 0.20% 0.76%
790m 370m 2.98% 2.29% 0.95% 1.79% 1.13% 1.90%
1.4b 130m 0.32% 0.18% 0.12% 0.18% 0.10% 0.57%
1.4b 370m 1.45% 1.07% 0.47% 0.90% 0.45% 1.19%
1.4b 790m 3.42% 3.00% 1.54% 3.04% 1.82% 2.87%
2.8b 130m 0.34% 0.26% 0.16% 0.25% 0.15% 0.53%
2.8b 370m 1.10% 0.80% 0.39% 0.66% 0.31% 0.87%
2.8b 790m 1.69% 1.31% 0.81% 1.71% 0.71% 1.56%
2.8b 1.4b 2.93% 2.58% 1.32% 2.67% 1.40% 2.30%
Table 8: The table lists the test results of Qwen (INT4 quantized), based on our proposed AR framework in fundamental language benchmarks.
LLM SLM WikiText2 WikiText103 text8 1BW Pile(part) LAMBADA
Qwen-int4 Qwen-int4 PPL $\downarrow$ PPL $\downarrow$ BPC $\downarrow$ PPL $\downarrow$ PLL $\downarrow$ PPL $\downarrow$
Single Model Baseline Performance
0.5b 13.62 16.22 1.04 46.70 12.04 41.77
1.5b 10.12 11.29 0.88 32.54 8.68 30.82
3b 9.00 9.75 0.82 28.33 7.70 27.82
7b 7.66 8.19 0.75 24.72 6.75 24.34
14b 6.53 6.96 0.67 21.79 6.03 21.28
Improvement of the evaluation metric relative to the best performing single model (%)
1.5b 0.5b 2.10% 2.10% 0.48% 1.54% 0.78% 1.53%
3b 0.5b 1.59% 1.50% 0.28% 1.18% 0.44% 1.16%
3b 1.5b 4.29% 4.39% 1.55% 4.16% 2.37% 3.90%
7b 0.5b 0.60% 0.60% 0.10% 0.43% 0.09% 0.53%
7b 1.5b 1.27% 1.38% 0.45% 1.54% 0.52% 1.32%
7b 3b 2.57% 2.84% 1.07% 3.01% 1.37% 2.28%
14b 0.5b 0.33% 0.46% 0.11% 0.23% 0.05% 0.24%
14b 1.5b 0.60% 0.88% 0.33% 0.69% 0.24% 0.49%
14b 3b 1.06% 1.38% 0.56% 1.17% 0.48% 0.73%
14b 7b 2.31% 2.99% 1.44% 2.48% 1.37% 1.59%
Pythia is an open-source suite of autoregressive Transformer models developed by EleutherAI to enable rigorous scientific research into large language model behavior and training dynamics, with fully public data provenance and checkpointing. It is a highly classic series of scaling law models. In our experiments, we used the 14M, 70M, 160M, 410M, 1B, 1.4B, and 2.8B variants.
Table 9: The table lists the test results of the classic multi-scale scaling law model Pythia based on our proposed AR framework in fundamental language benchmarks.
LLM SLM WikiText2 WikiText103 text8 1BW Pile(part) LAMBADA
Pythia Pythia PPL $\downarrow$ PPL $\downarrow$ PPL $\downarrow$ BPC $\downarrow$ PLL $\downarrow$ PPL $\downarrow$
Single Model Baseline Performance
14m 95.80 134.06 1.66 234.11 39.09 81.20
70m 39.16 52.15 1.42 100.97 19.97 49.40
160m 24.20 29.73 1.26 66.36 13.89 37.17
410m 16.14 18.23 1.08 47.20 9.84 28.55
1000m 13.71 15.03 1.02 40.08 8.59 25.72
1400m 12.48 13.39 0.97 35.32 7.97 24.06
2800m 10.92 11.54 0.92 31.27 7.23 22.12
Improvement of the evaluation metric relative to the best performing single model (%)
70m 14m 0.47% 0.38% 0.19% 0.24% 0.27% 1.18%
160m 14m 0.03% 0.01% 0.01% 0.02% 0.27% 0.39%
160m 70m 0.95% 0.51% 0.30% 0.81% 0.78% 1.33%
410m 14m 0.04% 0.04% 0.01% 0.11% 0.12% 0.41%
410m 70m 0.44% 0.14% 0.06% 0.56% 0.18% 0.66%
410m 160m 1.15% 0.50% 0.18% 1.38% 0.43% 1.19%
1b 14m 0.02% 0.02% 0.00% 0.01% 0.09% 0.35%
1b 70m 0.18% 0.05% 0.02% 0.17% 0.09% 0.46%
1b 160m 0.43% 0.15% 0.07% 0.44% 0.16% 0.69%
1b 410m 2.51% 1.91% 0.97% 1.91% 1.02% 1.93%
1.4b 14m 0.01% 0.02% 0.00% 0.00% 0.14% 0.28%
1.4b 70m 0.10% 0.04% 0.01% 0.05% 0.14% 0.38%
1.4b 160m 0.20% 0.09% 0.04% 0.16% 0.19% 0.54%
1.4b 410m 0.98% 0.66% 0.33% 0.65% 0.41% 0.99%
1.4b 1b 3.17% 2.66% 1.05% 2.24% 1.77% 2.40%
2.8b 14m 0.01% 0.02% 0.00% 0.00% 0.16% 0.20%
2.8b 70m 0.08% 2.42% 0.01% 0.02% 0.15% 0.26%
2.8b 160m 0.15% 0.06% 0.03% 0.07% 0.19% 0.39%
2.8b 410m 0.36% 0.27% 0.16% 0.30% 0.20% 0.53%
2.8b 1b 1.07% 0.85% 0.44% 0.82% 0.59% 1.08%
2.8b 1.4b 1.83% 1.68% 0.98% 1.95% 1.10% 1.75%
A series of experimental results based on models with different structures and purposes indicates that our proposed AR framework, which is based on relative overfitting, has achieved widespread effectiveness, and the results are in almost complete accordance with the AR law we proposed after thousands of experiments.
E Experimental Setting/Details
All experiments were conducted on an RTX 4090 graphics card.
E.1 Experimental Section 5.2
In Section 5.2, the custom models (SimpleTransformer) were trained using tokenizers derived from various reference pre-trained models (e.g., GPT2, Mamba).
The custom model employed a standard transformer encoder architecture with the following parameters:
  • Transformer Encoder Layers: 6
  • Attention Heads per Layer: 6
  • Hidden Size / Embedding Dimension: 384
[…]
  • Activation Function: GELU
  • Dropout Rate: 0.1
  • Maximum Sequence Length / Positional Embeddings: 1024
  • Vocabulary Size: Determined by the specific reference tokenizer used for each training run.
The training process for these models utilized the following hyperparameters:
  • Random Seed: 42
  • Batch Size (per device): 2
  • Gradient Accumulation Steps: 4 (Effective Batch Size: 8)
  • Optimizer: AdamW
  • Learning Rate: 5e-5
  • LR Scheduler: Cosine schedule with 10% warmup steps
  • Gradient Clipping: Maximum norm of 1.0
  • Mixed Precision Training: Enabled
E.2 Experimental Section 5.3
In Section 5.3, we describe the training of three types of sequential language models: LSTM, GRU, and Transformer.
For each architecture type, separate models were trained by varying the embedding dimension (embed_size) across the values: **16, 32, 128, 256, 512, 768, and 1024**.
The remaining architectural parameters were fixed based on the script’s defaults for each specific architecture type:
  • For LSTM and GRU models:
    • Hidden Size (hidden_size): 128
    • Number of Recurrent Layers (num_layers): 1
  • For Transformer models:
    • Number of Attention Heads (nhead): 2
    • Number of Transformer Encoder Layers (num_layers): 1
    • Sequence Length for Positional Encoding (seq_len): 35
  • Common to all architectures:
    • Dropout Rate: 0.2 (applied in relevant layers)
Note: The vocabulary size (vocab_size) was constructed based on the training data (WikiText2 train split), capped at a maximum of 10,000 tokens and requiring a minimum frequency threshold of 2.
The training process for all these models utilized the following hyperparameters:
  • Epochs: 5
  • Batch Size: 20
  • Optimizer: Adam
  • Learning Rate: 0.001
  • Loss Function: Cross-Entropy Loss (nn.CrossEntropyLoss)
F Licenses
  • GPT2 model: MIT license
  • Qwen models: Apache 2.0 license
  • RWKV models: Apache 2.0 license
  • Mamba models: Apache 2.0 license
  • Pythia models: Apache 2.0 license
  • ResNet models: Apache 2.0 license
  • ESM2 models: MIT license
  • WikiText2 dataset: CC BY-SA 3.0 license and GFDL
  • WikiText103 dataset: CC BY-SA 3.0 license and GFDL
  • text8 dataset: License not specified
  • 1BW (One Billion Word) dataset: Apache 2.0 license
  • Pile dataset: MIT license
  • LAMBADA benchmark: CC BY 4.0 license
  • CBT (Children’s Book Test) benchmark: GFDL
  • ARC benchmark: CC BY-SA 4.0 license